OpenStudy (anonymous):

Can some one solve this problem? Find the product: (3x-5)(x+3)

OpenStudy (mertsj):

Use FOIL

OpenStudy (anonymous):

Use FOIL. \[3x*x=?\] \[3x*3=?\] \[-5 * x = ?\] \[-5 * 3 = ?\] Then add up all the results.

OpenStudy (anonymous):

(3x-5)(x+3) take your 3x and multiply it to x and 3 individually.. 3x^2+3x+3 -5x-15 then combine like terms (i think )

OpenStudy (anonymous):

here are the answer? A. 3x^2+4x-15 B. 3x^-4x-15 C. 3x^2+14x-15 D. 3x^2-14x+15

OpenStudy (anonymous):

ok. well.. where i have the 3x+3 should be 9x so 3x^2+9x-5x-15 then combine like terms

OpenStudy (anonymous):

i might be wrong but im thinking its a

OpenStudy (anonymous):

ok thanxsz